OpenStudy (anonymous):

@jim_thompson5910 5. Graph the circle with center at (-1, 3) and radius 6. Label the center and at least four points on the circle. Write the equation of the circle.

jimthompson5910 (jim_thompson5910):

What do you get? This will use the same basic steps used last time.

OpenStudy (anonymous):

-1 and 3 and the radius of 6 but whats the equation?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well start with (x-h)^2 + (y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

the center is (h,k), so (h,k) = (-1,3) ----> h = -1 and k = 3

jimthompson5910 (jim_thompson5910):

the radius is 6, so r = 6

jimthompson5910 (jim_thompson5910):

what do you get when you plug all that in

OpenStudy (anonymous):

(x-(-1)^2+(y-3^2=6^2 now what

jimthompson5910 (jim_thompson5910):

(x-h)^2 + (y-k)^2 = r^2 (x-(-1))^2 + (y-3)^2 = 6^2 (x-(-1))^2 + (y-3)^2 = 36 (x+1)^2 + (y-3)^2 = 36

jimthompson5910 (jim_thompson5910):

So your equation is (x+1)^2 + (y-3)^2 = 36